x|x|=4 x =?

x|x|=4
x =?

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
x|x|=x^2,x>0
=-x^2 x<0
therefore x^2=4x x>0
x=0,4
-x^2=4x x<0
x=-4,0

thanks and regards
sunil kr
askIItian faculty